# From Age-Structured Trophic Networks to Applied Control : Stabilization and Harvesting Strategies for Non-Transitive Competition and the Dynamics of Mosquitoes

### v1

@@ 2.3. Case of a nonlinear nonlocal term
where each $f_{i}(Y(\cdot,t))$ is a given nonlinear functional of the age‑density $Y(\cdot,t)$. Using the condition (1.1), system (1.4) becomes
(2.24) $\left\{\begin{array}[]{ll}\partial_{t}Y(a,t)+\partial_{a}Y(a,t)+D(a)Y(a,t)+M(t).Y(a,t)=0&\text{ in }Q,\\ Y(0,t)=\displaystyle\int_{0}^{A}\bar{\beta}(a,P_{1}(t))Y(a,t)da,&\text{ in }Q_{T},\\ Y(a,0)=Y_{0}&\text{ in }Q_{A},\end{array}\right$.
[…]
@@
    (2.25) $Y=\left\{\begin{array}[]{ll}Y_{0}(a-t)e^{-\displaystyle\int_{a-t}^{a}(D(s)+M(t))ds}\qquad\text{for}\quad t\leq a,\\ \\ b(t-a)e^{-\displaystyle\int_{0}^{a}(D(s)+M(t))ds}\qquad\text{for}\quad t>a,\end{array}\right$.
@@
    (2.28) $F(t)=\displaystyle\int_{t}^{A}\bar{\beta}(a,P_{1})Y_{0}(a-t)e^{-\displaystyle\int_{a-t}^{a}(D(s)+M(s))ds}da,\qquad k(a,t)=\bar{\beta}(a,P_{1}(t))e^{-\displaystyle\int_{0}^{a}(D(s)+M(s))ds}$
@@
    (2.30) $|F(t)|\leq\|\bar{\beta}\|_{L^{\infty}}e^{-t(\bar{\mu+M)}}\|Y_{0}\|_{\mathcal{H}_{2}},\quad\text{and}\;|k(a,t)|\leq\|\bar{\beta}\|e^{-a(\bar{\mu}+M)}.$
@@
    Choosing $\lambda$ sufficiently large such that $\frac{\|\bar{\beta}\|}{\bar{\mu}+M+\lambda}<1$ ensures that $\mathcal{F}$ is a contraction on $L^{\infty}(0,T)^{N}$, and Banach’s fixed point theorem yields the existence and uniqueness of the solution $b(t)$. ∎
@@
Non-Kolmogorov case. Suppose the model is not of Kolmogorov type, i.e., the function $f$ does not take the form (1.2), which corresponds to a particular case. Then, the existence of solutions can be established by reformulating the problem as a Volterra integral equation. In particular, the function $b(t)=Y(0,t)$ satisfies
@@ Discussion
[…]
@@ 2.4. Steady states
The system (2.43) can be formulated either as a Kolmogorov-type system (when the functional responses used are of the Holling-Tanner type) or as another type of system (when using Holling Types II or III), and the analysis varies depending on this choice. In the case of a Kolmogorov-type system, we obtain a solution of the form $Y(a)=Y(0)\Pi_{1}(a)$ where $\Pi_{1}$ is defined in (2.50). We simplify the calculation by assuming that the newborns satisfy equation (1.5).
@@
(2.44) $\left\{\begin{array}[]{ll}\partial_{t}\widetilde{y}+\partial_{a}\widetilde{y}+D(a)\,\widetilde{y}\;+\;\tilde{y}f^{\prime}\bigl{(}\overline{Y}\bigr{)}\;=\;0&\text{in }Q,\\ \widetilde{y}(0,t)\;=\;\displaystyle\int_{0}^{A}\beta(a)\,\widetilde{y}(a,t)\,da&\text{in }Q_{T}.\end{array}\right$.
@@
(2.45) $\left\{\begin{array}[]{ll}\partial_{a}\,w\;+\;\bigl{(}\lambda I+D(a)\;+\;f^{\prime}\bigl{(}\overline{Y}(a)\bigr{)}\bigr{)}\,w\;=\;0&\text{in }\;(0,A),\\ w(0)\;=\;\displaystyle\int_{0}^{A}\beta(a)\,w(a)\,da\,.\end{array}\right$.
@@
Let us consider hypothesis (H4), $\bar{Y}$ be the stationary solution of (1.4). If $R_{0}<I_{\mathcal{H}^{2}},$ then the steady state of system(1.4) is locally asymptotically stable. Otherwise, if $R_{0}>I_{\mathcal{H}^{2}}$ then the steady state is unstable.
[…]
@@
Let us consider hypothesis (H4), $R_{0}<I_{\mathcal{H}^{2}}.$ Then, the system (1.4) is locally exponentially stable.
[…]
@@
    also generates a (exponentially stable) $C_{0}$‐semigroup on $\mathcal{H}_{2}$. ∎
@@ Discussion
[…]
@@ 3. Applications : Examples of models and stability analysis
(3.6) $[M\bullet_{r}x]_{i}:=\sum_{j=1}^{N}M_{ij}\,x_{j},\quad i=1,\dots,N.$
@@
that is, explicitly,
(3.8) $\big{[}\mathcal{B}(X(t))\big{]}_{i}=\sum_{j=1}^{N}a_{ij}\,\gamma_{ij}(t)\,X_{j}(t),\quad i=1,\dots,N.$
@@
(3.15) $\begin{cases}\partial_{t}X(a,t)+\partial_{a}X(a,t)=\bar{\Gamma}(t)X(a,t)-\big{(}M(a)+B\circ U(t)+\mathcal{B}(X(t))\big{)}\bullet_{r}\;X(a,t),&\text{in}\;Q,\\[6.0pt] X(0,t)=\displaystyle\int_{0}^{A}K(a)\,X(a,t)\,da,&\text{in}\;Q_{T},\\[6.0pt] X(a,0)=X_{0}(a),&\text{in}\;Q_{A}.\end{cases}$
@@
(3.17) $\begin{cases}\partial_{t}X(a,t)+\partial_{a}X(a,t)=-\big{(}M(a)+B\circ U(t)+\mathcal{B}(X(t)))\big{)}\;\bullet_{r}\;X(a,t),&\text{in}\;Q,\\[6.0pt] X(0,t)=\displaystyle\int_{0}^{A}K(a)\,X(a,t)\,da,&\text{in}\;Q_{T},\\[6.0pt] X(a,0)=X_{0}(a),&\text{in}\;Q_{A}.\end{cases}$
@@
[…]
    Thanks to Lemma (2.1), the operator
    (3.18) $\mathcal{A}_{l}:D(\mathcal{A}_{l})\subset\mathcal{H}_{2}\longrightarrow\mathcal{H}_{2},\quad\mathcal{A}_{l}\varphi=-\partial_{a}\varphi-M(a)\varphi,$
@@
    is the infinitesimal generator of a strongly continuous semi-group on $\mathcal{H}_{2}.$ Since the control $U$ is bounded, the result follows from Theorem 2.8. ∎
@@
In the study of stability for dynamic competition models, particularly in non-transitive settings, we establish a general stability result for non-transitive competition models by induction, using a reduced control localized on a single species. The approach treats first the three-species case, then the four-species case, to derive the inductive step that extends to the general $N+1$ species case. The single control, applied to one species, acts indirectly on the other system components.To synthesize the global feedback, we construct successive fictitious controls that partially stabilize each species; these controls are implemented step by step and recursively until the global control is obtained.
@@ 3.1. Three-species non-transitive competition
An age-structured system with three interacting populations can model cyclic competition observed in microbial communities, where each species produces a toxin that inhibits another. For example, Kerr et al. (2002) experimentally demonstrated a rock–paper–scissors dynamic using genetically modified Escherichia coli strains: species A kills B, B kills C, and C kills A. In the model, the nonlocal terms $\gamma_{ij}(t)$ describe the age-distributed inhibitory effects of one species on another. A control applied to species 1 may represent a targeted antibiotic treatment. This framework captures both experimentally validated and theoretically analyzed dynamics of cyclic dominance (Durrett and Levin, 1998; Frean et al., 2008; Nowak and Sigmund, 2004).
In this section we focus on the non-transitive competition model (3.20). Non-transitive competition is an ecological relationship in which species interactions do not follow a linear hierarchy but rather form a cycle. Concretely, species $x_{1}$ dominates $x_{3}$, $x_{3}$ dominates $x_{2}$, and $x_{2}$ in turn dominates $x_{1}$. This organization tends to promote stable diversity in the ecosystem by preventing any single species from achieving exclusive dominance. Moreover, such cycles foster the emergence of ecological niches and complex dynamics (oscillations, cyclic coexistence, etc.), phenomena commonly observed in natural communities.
@@ 3.1.1. Stability analysis
(3.21) $\begin{cases}\partial_{a}x^{*}_{1}(a)=-\left(\mu_{1}(a)+\zeta_{1}\right)x^{*}_{1}(a),&\text{ in }Q_{A},\\[8.0pt] \\ \partial_{a}x_{2}(a)=-\left(\mu_{2}(a)+\zeta_{2}\right)x^{*}_{2}(a),&\text{ in }Q_{A},\\[8.0pt] \\ \partial_{a}x_{3}(a)=-\left(\mu_{3}(a)+\zeta_{3}\right)x^{*}_{2}(a),&\text{ in }Q_{A},\\[8.0pt] x^{*}_{i}(0)=\displaystyle\int_{0}^{A}k_{i}(a)x_{i}^{*}(a)da,\quad i=1,...,3\end{cases}\qquad\text{with}\;\begin{cases}\zeta_{1}=\lambda_{2}+u^{*},\\ \zeta_{2}=\lambda_{3},\\ \zeta_{3}=\lambda_{1},\\ \lambda_{i}=\displaystyle\int_{0}^{A}g_{j}(a)x^{*}_{i}(a)da\end{cases}$
@@
(3.38) $\begin{cases}\mathcal{D}^{*}\pi_{0,1}(a)=\partial_{a}\pi_{0,1}(a)-\pi_{0,1}(a)(\mu_{1}(a)+\zeta_{1})+\pi_{0,1}(0)k_{1}(a),\quad\pi_{0,1}(A)=0,\\ \mathcal{D}^{*}\pi_{0,2}(a)=\partial_{a}\pi_{0,2}(a)-\pi_{0,2}(a)(\mu_{2}(a)+\zeta_{2})+\pi_{0,2}(0)k_{2}(a),\qquad\pi_{0,2}(A)=0,\\ \mathcal{D}^{*}\pi_{0,3}(a)=\partial_{a}\pi_{0,3}(a)-\pi_{0,3}(a)(\mu_{3}(a)+\zeta_{3})+\pi_{0,3}(0)k_{3}(a),\qquad\pi_{0,3}(A)=0,\end{cases}$
by following the strategy employed in [27]. Applying transformation (3.30) yields equation (3.34). ∎
@@ 3.1.1.1. Stability for $\psi_{i}\equiv 0$
(3.41) $V_{i}(\eta_{i})=\lambda_{i}\bigl{(}e^{\eta_{i}}-1-\eta_{i}\bigr{)},$
@@
Let us introduce the following notation:
$\phi_{i}(z_{j})=\lambda_{i}\bigl{(}e^{z_{j}}-1\bigr{)},\qquad\alpha_{i}(\eta_{j})=\eta_{j},\qquad\alpha_{i}(z_{j})=z_{j},$
@@
(3.44) $V_{3}(\eta_{2},z_{1},z_{2})=\theta\lambda_{2}\bigl{(}e^{\eta_{2}}-1-\eta_{2}\bigr{)}+\theta\lambda_{3}\big{(}e^{z_{1}}-1-z_{1}\big{)}+\lambda_{1}(e^{z_{2}}-1-z_{2}).$
@@
Under the proposed feedback law the system (3.40) is globally asymptotically stable. The feedback control constructed is uniformly bounded, though not necessarily nonnegative. Moreover, the control satisfies $u(t)>0$ for every $t>0,$ for every $\eta_{i}(0)$ belonging to the largest level set of $V_{3}(\eta_{2},z_{1},z_{2})$ within the set
(3.45) $\mathcal{K}=\left\{\eta\in\mathbb{R}^{3}\ \middle|\ \begin{aligned} &u^{*}+c_{3o}\phi_{1}(z_{2})-\theta\frac{1}{\lambda_{3}}\phi^{2}_{3}(z_{1})+(\dfrac{\lambda_{1}}{\lambda_{3}}-1-\theta)\phi_{3}(z_{1})+\frac{1}{\lambda_{3}}\phi_{3}(z_{1})\phi_{1}(z_{2})-\frac{1}{\lambda_{2}}\phi_{3}(z_{1})\phi_{2}(\eta_{2})\\ &-(\dfrac{\lambda_{3}}{\lambda_{2}}+1)\phi_{2}(\eta_{2})+\frac{\theta\phi_{3}(z_{1})}{\phi_{1}(z_{2})}\left(\frac{1}{\lambda_{2}}\phi_{3}(z_{1})\phi_{2}(\eta_{2})-\frac{1}{\lambda_{2}}\phi^{2}_{2}(\eta_{2})+(\dfrac{\lambda_{3}}{\lambda_{2}}-1)\phi_{2}(\eta_{2})+\,\phi_{3}(z_{1})\right)>0\end{aligned}\right\}.$
[…]
@@
    (3.46) $\begin{cases}\dot{\eta}_{2}=-\phi_{3}(\eta_{3}),\\ \dot{z}_{1}=-\phi_{1}(\eta_{1})-\dot{\eta}_{2},\\ \dot{z}_{2}=-\phi_{2}(\eta_{2})-\dot{z}_{1}+u^{*}-u\end{cases}$
@@
      To ensure decay, we choose the virtual control with $c_{1}>0$ such that
@@
      (3.53) $\dot{z}_{1}=-\phi_{1}(\eta_{1})+\,\lambda_{3}\big{(}e^{z_{1}+\eta_{2}}-1\big{)}.$
      We define the composite Lyapunov function for $(\eta_{2},z_{1})$ as
      (3.54) $\omega_{2}(\eta_{2},z_{1})=\omega_{1}(\eta_{2})+\lambda_{3}\big{(}e^{z_{1}}-1-z_{1}\big{)},$
@@
      (3.55) $\dot{\omega}_{2}(\eta_{2},z_{1})=-c_{1}\,\phi_{2}^{2}(\eta_{2})-\frac{c_{1}}{\lambda_{3}}\phi^{2}_{2}(\eta_{2})\,\phi_{3}(z_{1})-\phi_{2}(\eta_{2})\,\phi_{3}(z_{1})+\phi_{3}(z_{1})\,\dot{z}_{1}.$
@@
      (3.56) $\dot{\omega}_{2}(\eta_{2},z_{1})$ $=-c_{1}\,\phi_{2}^{2}(\eta_{2})-\phi_{3}(z_{1})\phi_{1}(\eta_{1})-\frac{c_{1}}{\lambda_{3}}\phi^{2}_{2}(\eta_{2})\,\phi_{3}(z_{1})-\phi_{2}(\eta_{2})\,\phi_{3}(z_{1})+\frac{c_{1}}{\lambda_{3}}\phi^{2}_{3}(z_{1})\phi_{2}(\eta_{2})+\phi^{2}_{3}(z_{1})+c_{1}\,\phi_{3}(z_{1})\,\phi_{2}(\eta_{2}).$
@@
      For the construction of the fictitious control $z_{1}$, we perform
      (3.57) $\phi_{1}\big{(}z_{1}\big{)}=c_{2}\,\phi_{3}(z_{1}),\quad\text{for}\;c_{2}>0.$

### v2

@@ 2.3. Case of a nonlinear nonlocal term
where each $f_{i}(Y(\cdot,t))$ is a given nonlinear functional of the age‑density $Y(\cdot,t)$. If the system is of Kolmogorov type, then model (1.4) takes the form
(2.24) $\left\{\begin{array}[]{ll}\partial_{t}Y(a,t)+\partial_{a}Y(a,t)+D(a)Y(a,t)+M(t).Y(a,t)=0&\text{ in }Q,\\ Y(0,t)=\displaystyle\int_{0}^{A}\bar{\beta}(a,P_{1}(t))Y(a,t)da,&\text{ in }Q_{T},\\ Y(a,0)=Y_{0}&\text{ in }Q_{A}.\end{array}\right$.
  • Proof of Theorem 2.8 :
@@
    (2.25) $Y=\left\{\begin{array}[]{ll}Y_{0}(a-t)e^{-\displaystyle\int_{a-t}^{a}(D(s)+M(t))ds}\hskip 18.49988pt\text{for}\qquad t\leq a,\\ b(t-a)e^{-\displaystyle\int_{0}^{a}(D(s)+M(t))ds}\hskip 18.49988pt\text{for}\qquad t>a,\end{array}\right$.
@@
    (2.28) $F(t)=\displaystyle\int_{t}^{A}\bar{\beta}(a,P_{1})Y_{0}(a-t)e^{-\displaystyle\int_{a-t}^{a}(D(s)+M(s))ds}da,\hskip 18.49988ptk(a,t)=\bar{\beta}(a,P_{1}(t))e^{-\displaystyle\int_{0}^{a}(D(s)+M(s))ds}$
@@
    (2.30) $|F(t)|\leq\|\bar{\beta}\|_{L^{\infty}}e^{-t(\bar{\mu+M)}}\|Y_{0}\|_{\mathcal{H}_{2}},\qquad\text{and}\;|k(a,t)|\leq\|\bar{\beta}\|e^{-a(\bar{\mu}+M)}.$
@@
    Choosing $\lambda$ sufficiently large such that $\frac{\|\bar{\beta}\|}{\bar{\mu}+M+\lambda}<1$ ensures that $\mathcal{F}$ is a contraction on $L^{\infty}(0,T)^{N}$, and Banach’s fixed point theorem yields the existence and uniqueness of the solution $b(t)$.
@@
Non-Kolmogorov case. Suppose the model is not of Kolmogorov type (as in G. F. Webb [33]), i.e., the function $f$ does not take the form (1.2), which corresponds to a particular case. Then, the existence of solutions can be established by reformulating the problem as a Volterra integral equation. In particular, the function $b(t)=Y(0,t)$ satisfies
@@ Discussion
Condition (2.41) is necessary and sufficient because it imposes a minimal initial density that ensures population persistence: it prevents extinction due to Allee effects and guarantees that ecological interactions (predation, resource availability, competition) remain sufficiently strong to sustain positive dynamics. Predator dietary flexibility and biodiversity enhance resilience by providing alternative food pathways and functional redundancy. Conversely, strict niche competition (Gause’s principle) may lead to exclusion. Importantly, ecosystem survival and stability (see [7]) depend on a combination of environmental and anthropogenic factors (land-use changes, human pressures, population management), which must be considered when interpreting or applying condition (2.41). Historically, Ronald Ross demonstrated the practical significance of such thresholds by showing that reducing mosquito density below a critical level can interrupt malaria transmission, an illustration of how threshold effects translate into actionable control strategies [4].
Mathematically, (2.41) reads as a constraint on the initial datum $Y_{0}(s)$ that guarantees existence, positivity and persistence of solutions to integro-differential equations; empirical cases of species recovery illustrate that well-functioning and properly managed ecosystems enable population rebound.
@@ 2.4. Steady states
The system (2.43) can be formulated either as a Kolmogorov-type system (when the functional responses used are of the Holling-Tanner type) or as another type of system (when using Holling Types II or III), and the analysis varies depending on this choice. In the case of a Kolmogorov-type system, we obtain a solution of the form $Y(a)=Y(0)\Pi_{1}(a)$ where $\Pi_{1}$ is defined in (2.50). We simplify the calculation by assuming that the newborns satisfy (H5i)(H5-i).
@@
(2.44) $\left\{\begin{array}[]{ll}\partial_{t}\widetilde{y}+\partial_{a}\widetilde{y}+D(a)\,\widetilde{y}\;+\;\tilde{y}f^{\prime}\bigl(\overline{Y}\bigr)\;=\;0\quad&\text{in }Q,\\ \widetilde{y}(0,t)\;=\;\displaystyle\int_{0}^{A}\beta(a)\,\widetilde{y}(a,t)\,da\,\quad&\text{in }Q_{T}.\end{array}\right$.
@@
(2.45) $\left\{\begin{array}[]{ll}\partial_{a}\,w\;+\;\bigl(\lambda I+D(a)\;+\;f^{\prime}\bigl(\overline{Y}(a)\bigr)\bigr)\,w\;=\;0\quad&\text{in }\;(0,A),\\ w(0)\;=\;\displaystyle\int_{0}^{A}\beta(a)\,w(a)\,da\,.\end{array}\right$.
@@
Let us consider assumption (H4), $\bar{Y}$ be the stationary solution of (1.4). If $R_{0}<I_{\mathcal{H}^{2}},$ then the steady state of system(1.4) is locally asymptotically stable. Otherwise, if $R_{0}>I_{\mathcal{H}^{2}}$ then the steady state is unstable.
  • Proof of Theorem 2.12 :
@@
Let us consider assumption (H4), $R_{0}<I_{\mathcal{H}^{2}}.$ Then, the system (1.4) is locally exponentially stable.
  • Proof of Theorem 2.13 :
@@
    also generates a (exponentially stable) $C_{0}$‐semigroup on $\mathcal{H}_{2}$.
@@ Discussion
For example, in a two-species predator–prey system one may have $R_{0,1}<1$ and $R_{0,2}<1$ without guaranteeing predator extinction or unrestrained prey outbreaks: intraspecific regulatory mechanisms, community functional redundancy, or predator generalism (the ability to switch to alternative prey) can stabilize the dynamics. In networks of five or more species, feedback loops (asymmetric competition, mutualism, cross-predation), reinforced by redundancy and trophic plasticity, often play a compensatory role: they preserve global stability despite heterogeneity in the $R_{0,i}$, allowing partial persistence, stable coexistence, or selective extinction. This illustrates that it is the topology and nature of interactions, rather than strict compliance with (2.52), that underlie ecosystem stability.
@@ 3. Applications : Examples of models and stability analysis
(3.6) $[M\bullet_{r}X]_{i}:=\sum_{j=1}^{N}M_{ij}\,x_{i},\quad i=1,\dots,N.$
@@
(3.14) $\begin{cases}\partial_{t}X(a,t)+\partial_{a}X(a,t)=-\mathcal{B}(X(t))-\big(M(a)+B\circ U(t)\big)\bullet_{r}\;X(a,t),&\text{in}\;Q,\\[6.0pt] X(0,t)=\displaystyle\int_{0}^{A}K(a)\,X(a,t)\,da,&\text{in}\;Q_{T},\\[6.0pt] X(a,0)=X_{0}(a),&\text{in}\;Q_{A}.\end{cases}$
@@
(3.16) $\begin{cases}\partial_{t}X(a,t)+\partial_{a}X(a,t)=-\mathcal{B}(X(t))-\big(M(a)+B\circ U(t)\big)\;\bullet_{r}\;X(a,t),&\text{in}\;Q,\\[6.0pt] X(0,t)=\displaystyle\int_{0}^{A}K(a)\,X(a,t)\,da,&\text{in}\;Q_{T},\\[6.0pt] X(a,0)=X_{0}(a),&\text{in}\;Q_{A}.\end{cases}$
@@
  • Proof of Proposition 3.1 :
    Thanks to Lemma 2.1, the operator
    (3.17) $\mathcal{A}_{l}:D(\mathcal{A}_{l})\subset\mathcal{H}_{2}\longrightarrow\mathcal{H}_{2},\qquad\mathcal{A}_{l}\varphi=-\partial_{a}\varphi-M(a)\varphi,$
@@
    is the infinitesimal generator of a strongly continuous semi-group on $\mathcal{H}_{2}.$ Since the control $U$ is bounded, the result follows from Theorem 2.8.
@@
In the study of stability for dynamic competition models, particularly in non-transitive settings, we establish a general stability result for non-transitive competition models by induction, using a reduced control localized on a single species. The approach treats first the three-species case, then the four-species case, to derive the inductive step that extends to the general $N$ species case. The single control, applied to one species, acts indirectly on the other system components.To synthesize the global feedback, we construct successive fictitious controls that partially stabilize each species; these controls are implemented step by step and recursively until the global control is obtained.
@@ 3.1. Three-species non-transitive competition
An age-structured system with three interacting populations can model cyclic competition observed in microbial communities, where each species produces a toxin that inhibits another. For example, Kerr et al. (2002) experimentally demonstrated a rock–paper–scissors dynamic using genetically modified Escherichia coli strains: species A kills B, B kills C, and C kills A. In the model, the nonlocal terms $\gamma_{ij}(t)$ describe the age-distributed inhibitory effects of one species on another. A control applied to species 1 may represent a targeted antibiotic treatment. This framework captures both experimentally validated and theoretically analyzed dynamics of cyclic dominance (Durrett and Levin, 1998; Frean et al., 2008; Nowak and Sigmund, 2004). A representative example of a non-transitive competition model is discussed in [21].
In this section, we focus on the non-transitive competition model (3.19). Non-transitive competition is an ecological relationship in which species interactions do not follow a linear hierarchy but rather form a cycle. Concretely, species $x_{1}$ dominates $x_{3}$, $x_{3}$ dominates $x_{2}$, and $x_{2}$ in turn dominates $x_{1}$. This organization tends to promote stable diversity in the ecosystem by preventing any single species from achieving exclusive dominance. Moreover, such cycles foster the emergence of ecological niches and complex dynamics (oscillations, cyclic coexistence, etc.), phenomena commonly observed in natural communities.
@@ 3.1.1. Stability analysis
(3.20) $\begin{cases}\partial_{a}x^{*}_{1}(a)=-\left(\mu_{1}(a)+\zeta_{1}\right)x^{*}_{1}(a),&\text{ in }Q_{A},\\[8.0pt] \\ \partial_{a}x_{2}(a)=-\left(\mu_{2}(a)+\zeta_{2}\right)x^{*}_{2}(a),&\text{ in }Q_{A},\\[8.0pt] \\ \partial_{a}x_{3}(a)=-\left(\mu_{3}(a)+\zeta_{3}\right)x^{*}_{2}(a),&\text{ in }Q_{A},\\[8.0pt] x^{*}_{i}(0)=\displaystyle\int_{0}^{A}k_{i}(a)x_{i}^{*}(a)da,\quad i=1,2,3\end{cases}\qquad\text{with}\;\begin{cases}\zeta_{1}=\lambda_{2}+u^{*},\\ \zeta_{2}=\lambda_{3},\\ \zeta_{3}=\lambda_{1},\\ \lambda_{i}=\displaystyle\int_{0}^{A}g_{j}(a)x^{*}_{i}(a)da\end{cases}$
@@
    (3.37) $\begin{cases}\mathcal{D}^{*}\pi_{0,1}(a)=\partial_{a}\pi_{0,1}(a)-\pi_{0,1}(a)(\mu_{1}(a)+\zeta_{1})+\pi_{0,1}(0)k_{1}(a),\qquad\pi_{0,1}(A)=0,\\ \mathcal{D}^{*}\pi_{0,2}(a)=\partial_{a}\pi_{0,2}(a)-\pi_{0,2}(a)(\mu_{2}(a)+\zeta_{2})+\pi_{0,2}(0)k_{2}(a),\hskip 18.49988pt\pi_{0,2}(A)=0,\\ \mathcal{D}^{*}\pi_{0,3}(a)=\partial_{a}\pi_{0,3}(a)-\pi_{0,3}(a)(\mu_{3}(a)+\zeta_{3})+\pi_{0,3}(0)k_{3}(a),\hskip 18.49988pt\pi_{0,3}(A)=0,\end{cases}$
    by following the strategy employed in [27]. Applying transformation (3.29) yields equation (3.33).
@@ 3.1.1.1. Stability for $\psi_{i}\equiv 0$
(3.40) $\omega(\eta_{i})=\lambda_{i}\bigl(e^{\eta_{i}}-1-\eta_{i}\bigr),$
@@
For clarity, we will denote by $z_{i}^{k}$ the $i$-th error variable of the system with $k$ species. Let us introduce the following notation:
$\phi_{i}(z^{k}_{j})=\lambda_{i}\bigl(e^{z^{k}_{j}}-1\bigr),\qquad\alpha_{i}(\eta_{j})=\eta_{j},\qquad\alpha_{i}(z^{k}_{j})=z^{k}_{j},$
@@
(3.43) $V_{3}(\eta_{2},z^{3}_{1},z^{3}_{2})=\theta\lambda_{2}\bigl(e^{\eta_{2}}-1-\eta_{2}\bigr)+\theta\lambda_{3}\big(e^{z^{3}_{1}}-1-z^{3}_{1}\big)+\lambda_{1}(e^{z^{3}_{2}}-1-z^{3}_{2}).$
@@
Under the proposed feedback law the system (3.39) is globally asymptotically stable. The feedback control constructed is uniformly bounded, though not necessarily nonnegative. Moreover, the control satisfies $u(t)>0$ for every $t>0,$ for every $\eta_{i}(0)$ belonging to the largest level set of $V_{3}(\eta_{2},z^{3}_{1},z^{3}_{2})$ within the set
(3.44) $\mathcal{K}=\left\{\eta\in\mathbb{R}^{3}\ \middle|\ \begin{aligned} &u^{*}+c_{3o}\phi_{1}(z^{3}_{2})-\theta\frac{1}{\lambda_{3}}\phi^{2}_{3}(z^{3}_{1})+(\dfrac{\lambda_{1}}{\lambda_{3}}-1-\theta)\phi_{3}(z^{3}_{1})+\frac{1}{\lambda_{3}}\phi_{3}(z^{3}_{1})\phi_{1}(z^{3}_{2})-\frac{1}{\lambda_{2}}\phi_{3}(z^{3}_{1})\phi_{2}(\eta_{2})\\ &-(\dfrac{\lambda_{3}}{\lambda_{2}}+1)\phi_{2}(\eta_{2})+\frac{\theta\phi_{3}(z^{3}_{1})}{\phi_{1}(z^{3}_{2})}\left(\frac{1}{\lambda_{2}}\phi_{3}(z^{3}_{1})\phi_{2}(\eta_{2})-\frac{1}{\lambda_{2}}\phi^{2}_{2}(\eta_{2})+(\dfrac{\lambda_{3}}{\lambda_{2}}-1)\phi_{2}(\eta_{2})+\,\phi_{3}(z^{3}_{1})\right)>0\end{aligned}\right\}.$
  • Proof of Theorem 3.7 :
@@
    (3.45) $\begin{cases}\dot{\eta}_{2}=-\phi_{3}(\eta_{3}),\\ \dot{z}^{3}_{1}=-\phi_{1}(\eta_{1})-\dot{\eta}_{2},\\ \dot{z}^{3}_{2}=-\phi_{2}(\eta_{2})-\dot{z}^{3}_{1}+u^{*}-u\end{cases}$
@@
      To ensure decay, we choose the fictitious control $\eta_{3}$ with $c_{1}>0$ such that
@@
      (3.52) $\dot{z}^{3}_{1}=-\phi_{1}(\eta_{1})+\,\lambda_{3}\big(e^{z^{3}_{1}+\eta_{2}}-1\big).$
      We define the composite Lyapunov function for $(\eta_{2},z^{3}_{1})$ as
      (3.53) $\omega_{2}(\eta_{2},z^{3}_{1})=\omega_{1}(\eta_{2})+\lambda_{3}\big(e^{z^{3}_{1}}-1-z^{3}_{1}\big),$
@@
      (3.54) $\dot{\omega}_{2}(\eta_{2},z^{3}_{1})=-c_{1}\,\phi_{2}^{2}(\eta_{2})-\frac{c_{1}}{\lambda_{3}}\phi^{2}_{2}(\eta_{2})\,\phi_{3}(z^{3}_{1})-\phi_{2}(\eta_{2})\,\phi_{3}(z^{3}_{1})+\phi_{3}(z^{3}_{1})\,\dot{z}^{3}_{1}.$
@@
      (3.55) $\dot{\omega}_{2}(\eta_{2},z^{3}_{1})$ $=-c_{1}\,\phi_{2}^{2}(\eta_{2})-\phi_{3}(z^{3}_{1})\phi_{1}(\eta_{1})-\frac{c_{1}}{\lambda_{3}}\phi^{2}_{2}(\eta_{2})\,\phi_{3}(z^{3}_{1})-\phi_{2}(\eta_{2})\,\phi_{3}(z^{3}_{1})+\frac{c_{1}}{\lambda_{3}}\phi^{2}_{3}(z^{3}_{1})\phi_{2}(\eta_{2})+\phi^{2}_{3}(z^{3}_{1})+c_{1}\,\phi_{3}(z^{3}_{1})\,\phi_{2}(\eta_{2}).$
@@
      To stabilize the dynamics $z^{3}_{1}$, we choose the constant $c_{2}$ such that
      (3.56) $\phi_{1}\big(z^{3}_{1}\big)=c_{2}\,\phi_{3}(z^{3}_{1}),\qquad\text{for}\;c_{2}>0.$